A coin of mass 10gram rolls along a horizontal lable with a velocity of 6 cm/s. What is  its kinetic energy?

Grade:12

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The net kinetic energy consists of both rotational and linear kinetic energy. Assuming the coin to be in the form of a disc, we have the net kinetic energy being given by:
K.E.=\frac{1}{2}mv^2+\frac{1}{2}I\omega^2=\frac{1}{2}mv^2+\frac{1}{2}\left ( \frac{1}{2}mr^2 \right )\omega^2=\frac{3}{4}mv^2=\frac{3}{4}\times\frac{10}{1000}\times\left (\frac{6}{100} \right )^2=0.000027 JRegards